Write in Logarithmic form
(4/11)^-3=1331/64

Answer by stanbon(48535)

Write in Logarithmic form
(4/11)^-3=1331/64
----------
Log means "exponent".
---------
The exponent you have is -3
The base is (4/11)
The result is 1331/64
------
Rewrite as
log(base 4/11) 1331/64 = -3
==========
Cheers,
Stan H.